Bond Prices:

  1. On Dec. 30th 2019, the interest rate (yield-to-maturity)on a ten-year Treasury bond was 1.827%. If the bond paid a 2% annual coupon for 10 years and had a par value of $1,000, what price would it have sold for on Dec. 30th?
  2. On Mar. 5nd 2020, the interest rate (yield-to-maturity) on a ten-year Treasury had dropped to 0.926%. Again, if a bond paid a 2% annual coupon for 10 years and had a par value of $1,000, what price would it have sold for on Mar. 5th?
Answer #1

a]

Price of a bond is the present value of its cash flows. The cash flows are the coupon payments and the face value receivable on maturity

Price of bond is calculated using PV function in Excel :

rate = 1.827% (YTM of bonds = market interest rate)

nper = 10 (Years remaining until maturity with 1 coupon payment each year)

pmt = 1000 * 2% (annual coupon payment = face value * coupon rate)

fv = 1000 (face value receivable on maturity)

PV is calculated to be $1,015.68

A1 X fi =PV(1.827%,10,1000*2%, 1000) B C D E F 4 A 1 ($1,015.68)

b]

Price of a bond is the present value of its cash flows. The cash flows are the coupon payments and the face value receivable on maturity

Price of bond is calculated using PV function in Excel :

rate = 0.926% (YTM of bonds = market interest rate)

nper = 10 (Years remaining until maturity with 1 coupon payment each year)

pmt = 1000 * 2% (annual coupon payment = face value * coupon rate)

fv = 1000 (face value receivable on maturity)

PV is calculated to be $1,102.13

A1 X foc =PV(0.926%,10,1000*2%, 1000) B C D E F A ($1,102.13)! 1